

********************************************************************
UARS WEEKLY STATUS REPORT - 05 Dec 2003
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
The PEM web site has moved. It has a new URL. Access the PEM web site
using http://wwwpem.space.swri.edu. The PEM data from the GUVI conjunction
periods has moved to http://wwwpem.space.swri.edu/guvi-pem.html.
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 2002365 has been
reprocessed and archived at the GSFC DAAC.
Level 2 PEM MEPS data file documentation is complete and submitted to
the DAAC. MEPS data files are version 1. Processed MEPS data files waiting
archive at the GSFC DAAC are between 1991274 and 1991329.
PEM continues to work with the GUVI team on the diagnosis of data from
the first conjunction period. Analysis of the two conjunction passes showed
that the aurora as detected by GUVI was at higher latitude than UARS.
GUVI images from other times from the first conjunction period show spatial
regions that both GUVI and PEM sampled when aurora was present. Separation
is several minutes to 10's of minutes when both instruments saw aurora.
PEM has seen no aurora during the second conjunction period. Due to a
spacecraft anomaly, no PEM data was collected during the third conjunction
period.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG attended the UARS Mission Status Review meeting and
the 3TDRSS Status meeting, both held on 2 December 2003.
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 22 November 2003 (Orbit 66743, GMT Day 326) through
28 November 2003 (Orbit 66833, GMT Day 332).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
NONE THIS REPORT PERIOD.
The beta angle ranged between the angles of 32.8 degrees to 34.7
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 11 December 2003 (DOY 345).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.95 to 1.29 degrees C. Additional
battery performance data for each battery on 21 November and 28
November 2003 are provided in Appendix D.
The clock error ranged from -10.1 to 09.8 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 02
SUSIM = 00
SOLSTICE = 00
PEM = 00
HALOE = 00
HRDI = 00
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 00
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The problem found with the CMS-B pollers has been resolved. Once the
CMS-B connectivity issue is resolved the remainder of the new CMS code
acceptance testing can be completed.
The FSW group has completed the testing of the new flight software
uplink and remove procedures. A few contingency tests will be run
next week and then the FSW portion of the 3 TDRS preparation will be
finished. End-to-end tests with all entities, DSN testing and FOD
training are the remaining tasks before the uplink can be executed.
MEETINGS
-None this report period.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 11 December 2003 (DOY 345).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 28 November 2003
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
66807 331/0538 03-305 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
66824 332/0854 03-306 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 28 November 2003
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 28 November 2003
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
66807 331/0538 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
66824 332/0858 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 28 November 2003
21 November 2003 - GMT Day 325 Beta = 31.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.2 0.000 0.0
2 +05.6/-156.8 +5.79/+4.50 +28.4/-14.4 25.1 1.040 20.9
3 +16.8/-72.8 +5.14/+3.87 +27.6/-15.2 25.0 1.030 20.7
28 November 2003 - GMT Day 332 Beta = 34.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.2 0.000 0.0
2 +11.2/-156.8 +5.79/+4.50 +27.2/-14.8 24.8 1.040 20.5
3 +16.8/-89.6 +4.82/+3.87 +26.8/-15.2 24.6 1.040 20.0
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 28 November 2003
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
03-054 ECLIPSE 66625 319/0421 N/A Configs Crashed
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
03-052 FOT See List Nov 03 30832 Generic Late Acquisition